# Invariant Graph Transformer for Out-of-Distribution Generalization

### v1

@@ 4.2. Performance Comparisons
The synthetic dataset GOOD-Motif serves as a crucial benchmark for evaluating structural shift resilience. On the basis split, GOODFormer improves classification accuracy by 11.0% over the strongest baselines, and by 14.4% on the size split. Furthermore, while other graph Transformer methods surpass the strongest general invariant learning methods and graph OOD methods on the size split due to architectural advantages, their OOD generalization ability on both the basis and size splits is significantly inferior compared to our GOODFormer. Our GOODFormer achieves the best performance on the synthetic dataset, demonstrating its superior capability to effectively handle structural shifts.
@@
To answer RQ2, we further conduct qualitative and quantitative evaluations to assess whether our proposed method can accurately identify invariant subgraphs. Specifically, we compare GOODFormer with baseline methods that also generate subgraphs, using the ground-truth invariant subgraphs as a reference. The results, presented in Figure 2, demonstrate that GOODFormer outperforms the baselines in capturing invariant subgraphs, laying the foundations for its efficacy to handle distribution shifts. Additional qualitative and quantitative results are provided in Appendix E.4.
@@ 4.4. Hyper-parameter Sensitivity
Figure 3. The impact of different hyper-parameters. Red lines denote the results of GOODFormer and grey dashed lines are the best results of all baselines.
To answer RQ4, we further evaluate the sensitivity of hyper-parameters, including the invariant learning hyper-parameter $\lambda$, the entropy loss hyper-parameter $\alpha_{E}$ and the encoding loss hyper-parameter $\alpha_{\mathrm{PSE}}$. For simplicity, we only report the results on the size split of GOOD-Motif in Figure 3, while results on other datasets show similar patterns and are provided in Appendix E.1.
The invariant learning hyper-parameter $\lambda$ is crucial. Setting $\lambda$ too high can lead to an excessive emphasis on invariance, causing other loss to be overshadowed. Conversely, a too small $\lambda$ may be insufficient to differentiate invariant and variant subgraphs. The entropy loss hyper-parameter $\alpha_{E}$ is also important. Setting $\alpha_{E}$ too high can cause the model to favor invariant parts with higher certainty, neglecting other components. Conversely, if $\alpha_{E}$ is too low, it fails to maintain a sharp softmax function. The hyper-parameter of the encoding loss $\alpha_{\mathrm{PSE}}$ moderately impacts model performance, suggesting the need for a careful balance between the encoding loss and other losses. Although an appropriate choice can further improve the performance, our method is not very sensitive to hyper-parameters and outperforms the best baselines with a wide range of hyper-parameters choices.
@@ 5.1. Graph Transformers
Transformers have seen tremendous success in various fields, including natural language processing (NLP) (Tunstall et al., 2022), computer vision (Han et al., 2022), and more recently, in the graph domain (Shehzad et al., 2024; Min et al., 2022). Compared to existing message-passing neural networks (MPNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Veličković et al., 2018), they show promise in addressing challenges like over-smoothing and over-squashing (Cai et al., 2023; Xing et al., 2024) as well as capturing long-term dependencies (Wu et al., 2021). Over time, various GTs approaches have emerged. Graphormer (Ying et al., 2021) was among the first, using a single dense attention mechanism while adding structural features through centrality and spatial encodings. SAN (Kreuzer et al., 2021) employed dual attention mechanisms on both fully-connected graphs and original graph edges, with Laplacian Positional Encodings for nodes. Subsequently, GraphTrans (Wu et al., 2021) marked the introduction of a hybrid architecture, initially stacking MPNN layers before fully connecting the graph. GraphGPS (Rampášek et al., 2022) provided a flexible framework for integrating message-passing networks with attention mechanisms, allowing for diverse combinations of PSEs. The field continues to evolve with recent GT variants such as Exphormer (Shirzad et al., 2023), Polynormer (Deng et al., 2024), NAGphormer (Chen et al., 2023), among others (Wu et al., 2024; Zhang et al., 2022a; Liu et al., 2024). However, all of these GTs rely on the assumption that the training and testing graph comes from the same distribution, leaving their effectiveness under distribution shifts unexplored. To our knowledge, we are the first to study distribution shifts of GTs.
@@ 6. Conclusion
In the paper, we introduce a novel GOODFormer designed to empower Graph Transformer’s capable of handling graph data distribution shifts. Three tailored modules, including entropy-guided invariant subgraph disentangler, evolving subgraph positional and structural encoder, and invariant GT learning, are jointly optimized to enable the graph Transformer to capture invariant relationships between predictive graph structural information and labels. Extensive experiments on synthetic and real-world datasets demonstrate the superiority of GOODFormer.
@@ Appendix B Limitations
Appendix C Theory
C.1. Assumption
Assumption C.1.
@@ C.2. Proofs
Following the DIR (Wu et al., 2022a), we can deduce that the oracle function $f_{Y}$ from Assumption C.1, which is given $\boldsymbol{\mathrm{G}}_{C}$, satisfies the principle. Moreover, there exists a unique structure equation model $f_{Y}$ that adheres to the principle. Consequently, if any model $h$ comprising Eq. (3) satisfies the principle, then this model effectively represents the oracle function $f_{Y}$ given $\boldsymbol{\mathrm{G}}_{C}$.
[…]
$0<\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{c}}\leq M,m\leq\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{s}}<0$
@@
[…]
@@
Appendix D Experimental Details
GOOD-Motif GOOD-HIV GOOD-SST2 GOOD-Twitter DrugOOD
basis size scaffold size length length assay
$\lambda$ 10 1 0.1 100 0.1 100 100
$\alpha_{E}$ 0.1 0.1 0.1 0.01 0.01 0.01 0.01
$\alpha_{\mathrm{PSE}}$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01
D.1. Datasets
We adopt five graph classification datasets to verify the effectiveness of GOODFormer. The dataset statistics are presented in Table 4.
Table 4. Statistics of datasets.
Dataset Number of Graphs
Train Validation Test
GOOD-Motif-basis/size 18,000 3,000 3,000
GOOD-HIV-scaffold 24,682 4,113 4,108
GOOD-HIV-size 26,169 4,112 3,961
GOOD-SST2-length 24,744 17,206 17,490
GOOD-Twitter-length 2,590 1,785 1,457
DrugOOD-assay 34,179 19,028 19,032
@@ D.1. Datasets
Methods Precision@10
DIR 0.203
±0.041
GSAT 0.396
±0.077
CIGA 0.247
±0.024
GOODFormer 0.698 ±0.069
D.2. Baselines
We provide detailed descriptions of baselines in our experiments as follows:
  • ERM: is trained with the standard empirical risk minimizing.
D.3. Additional Details of Optimization and Hyper-parameters
For $\mathrm{GT}^{M}(\cdot)$ employing the GraphGPS layer (Rampášek et al., 2022), along with MPNN $h_{\mathrm{PSE}}(\cdot)$ and $\mathrm{GT}^{\prime}(\cdot)$, the numer of the respective layers is 2, 3 and 1, respectively. For all GTs, the dimensionality of both graph-level and node-level representations is $d=128$, whereas for GNN-based baselines, $d=300$. The global pooling function employed is mean pooling. The hand-crafted PSE $z_{\mathrm{PSE}}$ is Laplacian eigenvectors encoding (LapPE). The hyper-parameter $\lambda$ in Eq. (15) is chosen from $\{10^{2},10^{1},10^{0},10^{-1}\}$. The hyper-parameter $\alpha_{E}$ in Eq. (17) is chosen from $\{10^{-1},10^{-2}\}$. The hyper-parameter $\alpha_{\mathrm{PSE}}$ and $\alpha_{S}$ in Eq. (17) are set at $10^{-2}$ and $10^{0}$, repectively. We report the mean results and standard deviations of three runs. The number of epochs for optimizing our proposed method and the GT baselines is set to 100, while other GNN-based baselines use the setting of the GOOD benchmark (Gui et al., 2022). We adopt Adaptive Moment Estimation (Adam) for optimization. The activation function is ReLU (Agarap, 2018). The selected hyper-parameters are reported in Table 3. Besides, our hyper-parameters have concrete meanings and do not necessitate a resource-intensive tuning:
  • Regarding $\alpha_{S}$: This parameter solely controls the influence of the classification loss for the classifier $w_{\tilde{s}}(\cdot)$. Actually, in our experiments, $\alpha_{S}$ is consistently set to 1.
Appendix E Additional Experimental Results
E.1. Hyper-parameter Sensitivity Analysis
We further evaluate the sensitivity of hyper-parameters, including the invariant learning hyper-parameter $\lambda$, the entropy loss hyper-parameter $\alpha_{E}$ and the encoding loss hyper-parameter $\alpha_{\mathrm{PSE}}$ on the basis and size splits of GOOD-Motif in Figure 4. Our method is not very sensitive to hyper-parameters and outperforms the best baselines with a wide range of hyper-parameters choices. Besides, many of our hyper-parameters have concrete meanings and do not necessitate a resource-intensive tuning:
  • Regarding $\alpha_{S}$: This parameter solely controls the influence of the classification loss for the classifier $w_{\tilde{s}}(\cdot)$. Actually, in our experiments, $\alpha_{S}$ is consistently set to 1.
E.2. Running Time Analysis
@@
E.3. Shift Level Insights
To evaluate GOODFormer’s OOD generalization ability in different shift levels, we add experiments with varying shift severity (controlled by b) on the SP-Motif Dataset (Wu et al., 2022a; Li et al., 2022c). The results are shown in Table 5. Our method consistently achieves the best performance.
E.4. Invariant Subgraph Discovering Results
To validate whether GOODFormer can effectively identify invariant subgraphs, we conduct a comparative study against subgraph-outputting baselines using ground-truth invariant subgraphs on the basis split of GOOD-Motif. Precision@10 is adopted as the evaluation metric, with results summarized in Table 7. Our analysis reveals that GOODFormer significantly outperforms baseline methods in invariant subgraph discovery, particularly under distribution shifts.

### v2

@@ 4.2. Performance Comparisons
The synthetic dataset GOOD-Motif serves as a crucial benchmark for evaluating structural shift resilience. On the basis split, GOODFormer improves classification accuracy by 11.0% over the strongest baselines, and by 7.8% on the size split. Furthermore, while other graph Transformer methods surpass most general invariant learning methods and graph OOD methods on the size split due to architectural advantages, their OOD generalization ability on both the basis and size splits is significantly inferior compared to our GOODFormer. Our GOODFormer achieves the best performance on the synthetic dataset, demonstrating its superior capability to effectively handle structural shifts.
@@
To answer RQ2, we further conduct a comparative study to evaluate whether our proposed method can accurately identify invariant subgraphs. Specifically, we compare GOODFormer with baseline methods that also explicitly generate subgraphs via edge masking, using the ground-truth invariant subgraphs as a reference on the basis split of the GOOD-Motif dataset. Both qualitative and quantitative evaluations are performed, with Precision@10 adopted as the evaluation metric. The results, summarized in Table 4 and visualized in Figure 2, demonstrate that GOODFormer significantly outperforms the baselines in capturing invariant subgraphs, laying the foundation for its efficacy in handling distribution shifts.
Methods Precision@10
DIR 0.203
±0.041
GSAT 0.396
±0.077
CIGA 0.247
±0.024
GOODFormer 0.698 ±0.069
@@ 4.4. Hyper-parameter Sensitivity
Figure 3. The impact of different hyper-parameters. Red lines denote the results of GOODFormer and grey dashed lines correspond to the results of EXPHORMER.
To answer RQ4, we evaluate the sensitivity of hyper-parameters, including the invariant learning hyper-parameter $\lambda$, the entropy loss hyper-parameter $\alpha_{E}$ and the encoding loss hyper-parameter $\alpha_{\mathrm{PSE}}$. For simplicity, we only report the results on the size split of GOOD-Motif in Figure 3, while results on other datasets show similar patterns and are provided in Appendix F.1.
The invariant learning hyper-parameter $\lambda$ is crucial. Setting $\lambda$ too high can lead to an excessive emphasis on invariance, causing other loss to be overshadowed. Conversely, a too small $\lambda$ may be insufficient to differentiate invariant and variant subgraphs. The entropy loss hyper-parameter $\alpha_{E}$ is also important. Setting $\alpha_{E}$ too high can cause the model to favor invariant parts with higher certainty, neglecting other components. Conversely, if $\alpha_{E}$ is too low, it fails to maintain a sharp softmax function. The hyper-parameter of the encoding loss $\alpha_{\mathrm{PSE}}$ moderately impacts model performance, suggesting the need for a careful balance between the encoding loss and other losses. Although an appropriate choice can further improve the performance, our method is not very sensitive to hyper-parameters and outperforms the best GT with a wide range of hyper-parameters choices.
@@ 5.1. Graph Transformers
Transformers have seen tremendous success in various fields, including natural language processing (NLP) (Tunstall et al., 2022), computer vision (Han et al., 2022), and more recently, in the graph domain (Shehzad et al., 2024; Min et al., 2022). Compared to existing message-passing neural networks (MPNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Veličković et al., 2018), they show promise in addressing challenges like over-smoothing and over-squashing (Cai et al., 2023; Xing et al., 2024) as well as capturing long-term dependencies (Wu et al., 2021). Over time, various GTs approaches have emerged. Graphormer (Ying et al., 2021) was among the first, using a single dense attention mechanism while adding structural features through centrality and spatial encodings. SAN (Kreuzer et al., 2021) employed dual attention mechanisms on both fully-connected graphs and original graph edges, with Laplacian Positional Encodings for nodes. Subsequently, GraphTrans (Wu et al., 2021) marked the introduction of a hybrid architecture, initially stacking MPNN layers before fully connecting the graph. GraphGPS (Rampášek et al., 2022) provided a flexible framework for integrating message-passing networks with attention mechanisms, allowing for diverse combinations of PSEs. The field continues to evolve with recent GT variants such as Exphormer (Shirzad et al., 2023), Polynormer (Deng et al., 2024), NAGphormer (Chen et al., 2023), among others (Wu et al., 2024b; Zhang et al., 2022a; Liu et al., 2024). However, all of these GTs rely on the assumption that the training and testing graph comes from the same distribution, leaving their effectiveness under distribution shifts unexplored. To our knowledge, we are the first to study distribution shifts of GTs.
@@ 6. Conclusion
In the paper, we introduce a novel GOODFormer designed to empower Graph Transformer’s capable of handling graph data distribution shifts. Three tailored modules, including entropy-guided invariant subgraph disentangler, evolving subgraph positional and structural encoder, and invariant GT learning, are jointly optimized to enable the graph Transformer to capture invariant relationships between predictive graph structural information and labels. Extensive experiments on synthetic and real-world datasets demonstrate the superiority of GOODFormer, along with detailed analyses and ablation studies. In the future, we plan to extend our method for more types of graphs, such as dynamic graphs (Wu et al., 2024a; Kazemi et al., 2020) and heterogeneous graphs (Hu et al., 2020; Zhang et al., 2019; Wang et al., 2019).
@@ Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (No. 62472018, 92570113) and the Fundamental Research Funds for the Central Universities JK2024-07.
@@ Appendix B Limitations
Appendix C Notations
Table 6. Notations.
Notation Description
$\mathbb{G},\mathbb{Y}$ The graph space and the label space
$\boldsymbol{\mathrm{G}},\boldsymbol{\mathrm{Y}}$ The random variables of graph and label
$G,Y$ An instance of graph and label
$\mathcal{G}$ A graph dataset
$\boldsymbol{\mathrm{G}}_{C}$ The invariant rationale (causal subgraph)
$\boldsymbol{\mathrm{G}}_{S}$ The variant subgraph (complement of $\boldsymbol{\mathrm{G}}_{C}$)
$f_{Y}$ The oracle structural equation model
$\tilde{\mathcal{P}}$ Any distribution over $\mathcal{P}(\boldsymbol{\mathrm{G}},\boldsymbol{\mathrm{Y}})$
$h$ The complete model comprising $h_{\hat{Y}}\circ h_{g}\circ\mathrm{Softmax}\circ h_{l}$
$h_{l}$ Layers preceding the softmax function in the disentangler $\Phi$
$h_{g}$ Layers following the softmax function in the disentangler $\Phi$
$h_{\hat{Y}}$ The predictor mapping $\boldsymbol{\mathrm{G}}_{C}$ to $\boldsymbol{\mathrm{Y}}$
$\hat{h}$ A function that satisfies the invariant principle
$n_{c}$ Number of distinct instances of $(i,j)_{c}$ (node relationships in $\boldsymbol{\mathrm{G}}_{C}$)
$n_{s}$ Number of distinct instances of $(i,j)_{s}$ (node relationships in $\boldsymbol{\mathrm{G}}_{S}$)
$M$ Upper bound of $\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{c}}$
$m$ Lower bound of $\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{s}}$
$t$ Temperature parameter
$H$ Entropy
$p_{i}$ Probability of the $i$-th class
$e_{i}$ Logit value for the $i$-th class
$\Phi$ The subgraph disentangler
We summarize the key notations and the corresponding descriptions in Table 6.
Appendix D Theory
D.1. Assumption
Assumption D.1.
@@ D.2. Proofs
Following the DIR (Wu et al., 2022a), we can deduce that the oracle function $f_{Y}$ from Assumption D.1, which is given $\boldsymbol{\mathrm{G}}_{C}$, satisfies the principle. Moreover, there exists a unique structure equation model $f_{Y}$ that adheres to the principle. Consequently, if any model $h$ comprising Eq. (3) satisfies the principle, then this model effectively represents the oracle function $f_{Y}$ given $\boldsymbol{\mathrm{G}}_{C}$.
For simplicity, we denote $h=h_{\hat{Y}}\circ h_{g}\circ\mathrm{Softmax}\circ h_{l}$, where $h_{l}$ is the layers preceding the softmax function, and $h_{g}$ is the layers following the softmax function in the distangler $\Phi$. let’s denote $\hat{h}=\hat{h}_{\hat{Y}}\circ\hat{h}_{g}\circ\mathrm{Softmax}\circ\hat{h}_{l}$ as a function that satisfies the principle. By definition, $\hat{h}_{g}(\mathrm{Softmax}(\hat{h}_{l}(\boldsymbol{\mathrm{G}})))=\boldsymbol{\mathrm{G}}_{C}$ and $\hat{h}_{\hat{Y}}(\boldsymbol{\mathrm{G}}_{C})=\boldsymbol{\mathrm{Y}}$.
Due to two key reasons:
  • $\hat{h}_{l}$ is the sole component responsible for distinguishing $\boldsymbol{\mathrm{G}}_{C}$ and $\boldsymbol{\mathrm{G}}_{S}$ from $\boldsymbol{\mathrm{G}}$;
  • $\hat{h}_{l}$ includes constraints (e.g., BatchNorm) that bound feature vectors and attention,
$\hat{h}_{l}(\boldsymbol{\mathrm{G}})$ is guaranteed to stay within a controlled range. Thus, for all nodes $(i,j)$:
$0<\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{c}}\leq M,m\leq\hat{h}_{l}(\boldsymbol{\mathrm{G}})_{(i,j)_{s}}<0,$
@@
To prove that for any $i$ and $\epsilon>0$, there must exist a datum $(G,Y)$ such that $(\mathrm{Softmax}(\hat{h}_{l}(G)))_{(i,j)_{c}}-(\mathrm{Softmax}(\hat{h}(G)))_{(i,j)_{s}}<\epsilon$ for all $(i,j)_{c}$ and $(i,j)_{s}$, we denote $e_{c}=\hat{h}_{l}(G)_{(i,j)_{c}}$ and $e_{s}=\hat{h}_{l}(G)_{(i,j)_{s}}$. Assume that there are $n_{c}$ distinct instances of $(i,j)_{c}$ and $n_{s}$ distinct instances of $(i,j)_{s}$. The expression simplifies as follows:
@@
Appendix E Experimental Details
E.1. Datasets
We adopt five graph classification datasets to verify the effectiveness of GOODFormer.
@@ E.1. Datasets
E.2. Baselines
We provide detailed descriptions of baselines as follows:
  • ERM: is trained with the standard empirical risk minimizing.
  • GREA (Liu et al., 2022): proposes an efficient graph rationalization framework that leverages environment-based augmentations to create virtual training examples. By performing all operations in a latent space, it avoids the computational overhead of explicit graph decoding and encoding while improving rationale identification accuracy.
  • CAL+ (Sui et al., 2024): introduces a causal learning framework that addresses OOD generalization in graph classification by mitigating the confounding effect of shortcut features. It leverages backdoor adjustment through a memory bank to combine causal features with diverse shortcuts, while using prototypes to ensure intra-class causal consistency.
Appendix F Additional Experimental Results
F.1. Hyper-parameter Sensitivity Analysis
We further evaluate the sensitivity of hyper-parameters, including the invariant learning hyper-parameter $\lambda$, the entropy loss hyper-parameter $\alpha_{E}$ and the encoding loss hyper-parameter $\alpha_{\mathrm{PSE}}$ on the basis and size splits of GOOD-Motif in Figure 4. Our method is not very sensitive to hyper-parameters and outperforms both the general OOD generalization method CIGA and graph Transformer EXPHORMER with a wide range of hyper-parameters choices. Besides, many of our hyper-parameters have concrete meanings and do not necessitate a resource-intensive tuning:
[…]
F.2. Running Time Analysis
@@
F.3. Shift Level Insights
To evaluate GOODFormer’s OOD generalization ability in different shift levels, we add experiments with varying shift severity (controlled by b) on the SP-Motif Dataset (Wu et al., 2022a; Li et al., 2022c). The results are shown in Table 7. Experimentally, GOODFormer performs strongly under moderate distribution shifts, though some degradation occurs under extreme shifts, mainly due to decreased accuracy in invariant subgraph discovery. Performance may also be limited in more extreme scenarios-such as semantic shifts in OOD detection (Cai et al., 2025) with unseen invariant subgraphs. That said, we wish to emphasize that GOODFormer is specifically designed for OOD generalization.
F.4. Graph Transformer vs. MLP as Predictor
We employ GTs as predictors instead of MLPs to enhance node representations with explicit structural information. To validate whether GTs indeed outperform standard MLPs in this role, we conduct a comparative ablation study, substituting the GT in Eq. (13) with an MLP, with results summarized in Table 9. Our analysis confirms that GTs consistently achieve superior performance compared to general MLPs.
Table 9. Performance comparison between GT and MLP predictors.
GOOD-Motif GOOD-HIV
basis size scaffold size
MLP 70.66
±9.36
79.36
±5.16
68.34
±4.43
59.70
±3.06
GT 81.97 ±16.74 80.02 ±7.94 70.38 ±0.27 62.69 ±0.11